# Zero-shot Prompt-based Video Encoder for Surgical Gesture Recognition

### v1

@@ Abstract
Purpose: Surgical video is an important data stream for gesture recognition. Thus, robust visual encoders for those data-streams is similarly important.
Methods: Leveraging the Bridge-Prompt framework, we fine-tune a pre-trained vision-text model (CLIP) for gesture recognition in surgical videos. This can utilize extensive outside video data such as text, but also make use of label meta-data and weakly supervised contrastive losses.
@@
Conclusion Bridge-Prompt and similar pre-trained+fine-tuned video encoder models present significant visual representation for surgical robotics, especially in gesture recognition tasks. Given the diverse range of surgical tasks (gestures), the ability of these models to zero-shot transfer without the need for any task (gesture) specific retraining makes them invaluable.
@@ 1 Introduction
Proposed intra-operative robotic gesture recognition systems are described as enabling real-time feedback, assisting in multi-arm guidance, and facilitating automated sub-tasks. However, current methods [1, 2, 3] for gesture recognition attempt to estimate a fixed set of supervised gestures from contemporaneously collected video and/or kinematic datastreams [4, 5, 6, 7]. These systems are often trained in a fully supervised manner, which requires frame-wise gesture labels for the entire training set, and importantly require specification of the target labels during the whole training phase.
In the present work, we investigate using weak supervision and text augmentation for pre-training a surgical gesture recognition model. We focus on improving visual feature extraction from video data streams, as these data offer rich information about surgical gestures (surgemes), and kinematic data require access to the research API, limiting current use cases. We use the Bridge-Prompt [8] framework which specifically relaxes the fully supervised constraints: Bridge-Prompt is able to use large weakly supervised datasets, which are relatively inexpensive and numerous in comparison to densely annotated data, and our experiments suggest that Bridge-Prompt generalizes well to unseen labels outside of pre-training phase (i.e., novel gestures).
We believe that Bridge-Prompt and similar “pre-trained+fine-tuned” image encoder models will be critical for any system implementing gesture recognition, and for other visual-understanding tasks in surgical robotics. As procedures vary greatly in gesture vocabulary, a trained recognition system will need the capacity to generalize without extensive retraining; that is to say that it will need zero-shot capability. We believe that sufficiently expressive visual features will allow for strong performance from simple prediction methods with minimal training. Constructing a good feature set will enable generalization to related but unseen tasks.
In order to validate these claims and evaluate Bridge-Prompt’s overall efficacy and zero-shot capability, we demonstrate usage on the JIGSAWS [9] and RARP-45 [5] dataset, simulating both standard and zero-shot use cases on cross-validated training schemes. We fine-tune an encoder contrastively on the training fold data, varying experimentally the amount of label information (both annotation presence/absence and label text). We then compare the performance of experimental cases and standard baselines from the literature using a standard prediction recognition model (MS-TCN++, [10]). Our experiments show the benefit of using Bridge-Prompt image encoders for gesture recognition tasks.
@@
  • We provide open-source code [^1] https://github.com/mx-ethan-rao/Robotic_Gesture_Tools for pre-training encoders with Bridge-Prompt
@@ 2 Related Work
Surgical robotic gesture recognition has an extensive history of study [11, 12, 13]. While multiple disparate modalities are each reasonable to incorporate into a recognition system, video is by far the most popular [6, 4, 14, 15], followed by integration of video with robotic kinematic data [16, 17, 18, 5], and discrete surgical event streams [19]. Notably, almost all proposed methods rely on video data, or on video derived features such as optical flow [20, 21].
Gesture recognition is usually considered in a temporal context, and thus many of the prediction models are taken from time-series prediction and forecasting domains: HMM, LSTMs, temporal convolution, and attention methods [22, 23, 24, 5]. The Temporal Convolutional Network (TCN)[24] has emerged as a common selection for the majority of deep learning based temporal gesture studies [25, 10]. Our objective is to test the effectiveness of differing image encoders, thus we use a generic TCN implementation throughout the experiments.
@@ 3 Method
This section describes the two parts of our model: 1) the video encoder (Bridge-Prompt) and 2) the downstream gesture recognition model. The former we construct during pre-training and is the focus of our empirical study, while the latter is used in evaluation of each of the encoders, including both the proposed encoder (Bridge-Prompt) and baseline encoders. Though neither is first proposed by this manuscript, we include descriptions of both, as understanding their nuances (or simplicity, in the case of the downstream model) is necessary for contextualizing experiments.
Bridge-Prompt Pre-training Architecture
@@
To measure the effectiveness of prompt-based video encoder, we freeze the weights in the video encoder and do frame-wise visual embedding. Next, we train a predictive model for gesture recognition based on the visual embedding. We choose to use an MS-TCN++[10] as our downstream gesture recognition model. Temporal Convolution Networks have become a common choice in action segmentation, and the MS-TCN++ is a refinement of the original TCN. MS-TCN++ uses a simple two-stage training and a slightly modified convolutional configuration with dilations. We avoid deeper or more nuanced architectures (such as those incorporating attention, or longer context windows) to ensure that performance is due to the quality of the features and not the complexity of the classifier itself.
@@ 4 Experiments
Implementation: We implement multiple Bridge-Prompt variants as well as two baseline image encoders (3DResNet [3] and I3D [1]) in PyTorch. Each Bridge-Prompt implementation is composed of a backbone image encoder (either the default ViT-B/16 [34] or ResNet-50 [35]) and the GPT-2 text encoder [33] which is discarded after training. These backbones are initialized using standard pre-trained weight-sets [^3] Both Bridge-Prompt and 3DResNet methods prescribe the use of weights from another task and training phase before the pre-training phase; in this paper “pre-trained” weights are from this first phase, while “pre-training” are the actions performed on gesture recognition data.. The Bridge-Prompt encoders are then fine-tuned using the Adam optimizer minimizing the sum of the losses in Eq. 3 in what we are calling the pre-training phase.
After pre-training we freeze the weights of each image encoder, then train a standard prediction network for the supervised task (MS-TCN++ [10]). This second directly supervised phase we call the training phase. The training phase has its own hold-out test set and it is on this hold-out that we report performance metrics. For JIGSAWS we use Leave-One-User-Out (LOUO) [9] cross-validation, and for RARP-45 we choose 10 videos (out of 36 total) as the test set.
@@
Training time: All experiments were run on NVIDIA A40 GPUs or NVIDIA A5000 GPUs. The pre-training phase for all Bridge-Prompt variants was conducted on two GPUs, otherwise only a single GPU was used. In JIGSAWS, pre-training for 50 epochs for each tested variant of Bridge-Prompt takes approximately 8 hours using two A40 GPUs. It takes 5 minutes to train the MS-TCN++ during the supervised training phase, using pre-extracted image encodings.
Performance metrics: We assess the outcomes using five standard evaluation metrics, Accuracy, Edit Distance, and F1@{10,25,50} [24]. For JIGSAWS we condition this on pre-training task (Knot Tying or Suturing), and for zero-shot cases we also condition on each unseen label.
@@
We first measure the quality of the features learned by the various encoders operating under normal (non-zero-shot) conditions. Results are reported in Table 1 and 4. These results show that Bridge-Prompt improves gesture recognition performance as measured by all but one performance metric. Further, due to the similarity between Bridge-Prompt performance with either ResNet50 or ViT backbones, this performance gain does not appear to be due to the transformer architecture of ViT. However, ResNet50 backbones appear to be slightly less stable in RARP-45 training; across multiple runs we experienced exploding gradients, and the method failed to converge. Tuning the learning-rate might resolve this issue, but we did not have the resources to tune that parameter. We include JIGSAWS results reported in van Amsterdam et al. 2022 for contextualization [^4] Van Amsterdam et al. 2022 [5] reports results for models that include input data from both video and robotic kinematic streams, and thus are not entirely comparable for selecting video encoders. Moreover, the particular architecture, training details, and data split cannot be reproduced without access to their codebase. We exclude their RARP-45 reported values as they evaluate on all 45 videos, but only 36 videos in RARP-45 are publicly available., as we believe this to be a state-of-the-art contemporary system utilizing all available data streams (both visual and kinematic), and (presumably) many model architecture optimizations and procedure refinements.
Table 1: Here we present gesture recognition results on the JIGSAWS datatset for 3DResNet, I3D, and Bridge Prompt (BP) with both ResNet50 and ViT variants, as well as the Bridge Prompt architecture with no RARP-45 specific pre-training (CLIP). We provide the MA-TCN[C] numbers reported by van Amsterdam et al. [5] for overall context, but MA-TCN[C] uses both the video and kinematic data streams and a more complex prediction model.
@@
Zero-shot Capability: We measure the efficacy of the Bridge-Prompt image encoder for describing unseen gestures, i.e., zero-shot generalization capability. This is done by selectively training encoders using only subsets of the gestures: for JIGSAWS, we test subsets with only gestures 1-5, only gestures 1-10, and no gestures at all (the “pure CLIP [27] encoder”), presenting results in Table 7. We disaggregate performance by task, gesture, and model in Table 3. The 1-10 gesture model necessarily contains 6 and 8, and thus performance is not reported. For RARP-45 we test subsets with only gestures 1-3 and gestures 1-6, presenting results in Table 4. Per-gesture performance disaggregation is provided as a table in the Appendix. For JIGSAWS, we also trained video encoders on one task but evaluated them on the other (“cross-task”), reported at the bottom of Table 7. This experiment is not possible in RARP-45, as it only has one task. In general we find that Bridge-Prompt pre-training with only a subset of gestures or even on a different task still provides improvement for most performance measures. This, to us, indicates that Bridge-Prompt has relatively high zero-shot capability.
Table 2: Above the double line, we show zero-shot gesture recognition on JIGSAWS with different sets of pre-trained labels (%), along with an All gestures supervised case for reference, and encoders using indices (e.g., “Gesture 2”) instead of text descriptions during pre-training (“ind”). “No gestures” has no pre-training. Below the double line, we show cross-task zero-shot generalization, using the Knot-Tying pre-trained model for Suturing supervised training/evaluation, and vice versa.
@@
Table 3: We show zero shot F1@10 scores for single gestures with different sizes of pre-training sets (%) that exclude those gestures, along with an All gestures supervised case for reference, and encoders using indices (e.g., “Gesture 2”) instead of text descriptions during pre-training (“ind”).
@@
Table 4: Here we present gesture recognition results on the RARP-45 datatset for 3DResNet, I3D, and Bridge Prompt (BP) with the ViT variant, as well as a “no RARP-45 specific pre-training” case (CLIP). BP-ResNet50 failed to converge. Above the double line, we show the standard (non-zero-shot) performance. Below the double line, we show zero-shot surgical gesture recognition with different sets of pre-trained labels (%), and the same encoders using indices (e.g., “Gesture 2”) instead of text descriptions during pre-training (“ind”).
Acc. Edit F1@10 F1@25 F1@50
3DResNet 66.97 76.95 71.76 66.34 52.90
I3D 65.95 74.41 71.52 65.35 51.23
CLIP-ViT 70.00 73.88 74.94 73.30 59.95
CLIP-ResNet50 68.93 71.85 70.83 66.90 55.09
BP-ResNet50 - - - - -
BP-ViT 77.36 72.16 76.29 74.50 65.32
3 gestures 77.78 75.09 80.96 79.28 70.60
3 gestures (ind) 78.93 75.80 80.82 79.38 69.78
6 gestures 79.19 71.85 76.60 73.68 65.17
6 gestures (ind) 79.73 76.51 80.70 79.06 69.01
[…]
Text description ablation: Finally, we show evidence that text descriptions provide negligible benefit for Bridge-Prompt encodings. We measured the value of gesture text descriptions by ablating them to simple “Gesture Index” categorical descriptors. For example, gesture 9 could be described as “using right hand to help tighten suture” or as “Gesture 9”. Results from these experiments are included in Table 7, 3, and 4.

### v2

@@ Abstract
Purpose: In order to produce a surgical gesture recognition system that can support a wide variety of procedures, either a very large annotated dataset must be acquired, or fitted models must generalize to new labels (so called “zero-shot” capability). In this paper we investigate the feasibility of latter option.
Methods: Leveraging the Bridge-Prompt framework, we prompt-tune a pre-trained vision-text model (CLIP) for gesture recognition in surgical videos. This can utilize extensive outside video data such as text, but also make use of label meta-data and weakly supervised contrastive losses.
@@
Conclusion Bridge-Prompt and similar pre-trained+prompt-tuned video encoder models present significant visual representation for surgical robotics, especially in gesture recognition tasks. Given the diverse range of surgical tasks (gestures), the ability of these models to zero-shot transfer without the need for any task (gesture) specific retraining makes them invaluable.
@@ 1 Introduction
Proposed intra-operative robotic gesture recognition systems are described as enabling automation, rapid skill assessment and pedagogic feedback, and more general surgical support [1]. However, current methods [2, 3, 4] for gesture recognition attempt to estimate a fixed set of supervised gestures from contemporaneously collected video and/or kinematic datastreams [5, 6, 7, 8]. These systems are often trained in a fully supervised manner, which requires frame-wise gesture labels for the entire training set, and importantly require specification of the target labels during the whole training phase.
We believe that in order to produce a system that can provide support to wide variety of procedures, gesture recognition as a sub-field will either need a) a large number of annotated datasets for each procedure as previous fully supervised methods demand, or b) a system that can generalize well to new label sets with limited supervision. The latter option is likely more efficient given the expense of annotation, and fits within the “zero-shot learning” paradigm, where a general representation is reused for labelling tasks unseen during training [9].
We thus investigate the feasibility of such a zero-shot gesture recognition method using weak supervision and text augmentation. We focus on improving visual feature extraction from video data streams, as these data offer rich information about surgical gestures (surgemes) and kinematic data require access to the research API, limiting current use cases. We use the Bridge-Prompt [10] framework which specifically relaxes the fully supervised constraints: Bridge-Prompt is able to use large weakly supervised datasets, which are relatively inexpensive and numerous in comparison to densely annotated data, and our experiments suggest that Bridge-Prompt generalizes well to unseen labels (i.e., novel gestures).
In order to validate these claims and evaluate Bridge-Prompt’s overall efficacy and zero-shot capability, we demonstrate usage on the JIGSAWS [11] and RARP-45 [6] dataset, simulating both standard and zero-shot use cases on cross-validated training schemes. We compare image encoders with differing configurations, varying experimentally the amount of label information (both annotation presence/absence and label text). We then compare the performance of experimental cases and standard baselines from the literature using a standard prediction recognition model (MS-TCN++, [12]). Our experiments show the benefit of using Bridge-Prompt image encoders for gesture recognition tasks.
@@
  • We provide open-source code [^1] https://github.com/mx-ethan-rao/Robotic_Gesture_Tools for prompt-tuning encoders with Bridge-Prompt
Terminology
Throughout this manuscript we use the phrase “pre-trained” exclusively to denote previously trained instances of encoder networks that we either directly use without modifying or further train using a different weakly-supervised loss function. We refer to this further training phase as “prompt tuning”.
No matter if an encoder is only pre-trained or incorporates prompt-training, its weights are then frozen and used unchanged for the supervised training phase, wherein a supervised network will be trained to predict gesture labels from the frozen embeddings.
@@ 2 Related Work
Surgical robotic gesture recognition has an extensive history of study [13, 14, 15]. While multiple disparate modalities are each reasonable to incorporate into a recognition system, video is by far the most popular [7, 5, 16, 17], followed by integration of video with robotic kinematic data [18, 19, 6], and discrete surgical event streams [20]. Notably, almost all proposed methods rely on video data, or on video derived features such as optical flow [21].
Gesture recognition is usually considered in a temporal context, and thus many of the prediction models are taken from time-series prediction and forecasting domains: HMM, LSTMs, temporal convolution, and attention methods [22, 23, 24, 6]. The Temporal Convolutional Network (TCN) has emerged as a common selection for the majority of deep learning based temporal gesture studies [24, 25, 12]. MS-TCN and MS-TCN++ [25, 12] are the most relevant instances of this family of temporal modeling methods, as they were specifically designed for temporal action segmentation, and have been used in several surgical workflow analysis tasks [26, 27]. Our objective is to test the effectiveness of differing image encoders and not the prediction head architectures, thus we use a generic MS-TCN++ implementation throughout the experiments.
@@ 3 Method
This section describes the two parts of our model: 1) the video encoder (Bridge-Prompt) and 2) the downstream gesture recognition model. The former we construct during prompt-tuning and is the focus of our empirical study, while the latter is used in evaluation of each of the encoders, including both the proposed encoder (Bridge-Prompt) and baseline encoders. Though neither is first proposed by this manuscript, we include descriptions of both, as understanding their nuances (or simplicity, in the case of the downstream model) is necessary for contextualizing experiments.
Bridge-Prompt Prompt-tuning Architecture
@@
To measure the effectiveness of prompt-based video encoder, during the evaluation we freeze the weights in the video encoder and do frame-wise visual embedding. The frame encoder is fixed, and we train a predictive model for gesture recognition based on that fixed visual embedding. Our chosen downstream predictive model is the MS-TCN++[12]Temporal Convolution Networks have become a common choice in action segmentation, and the MS-TCN++ is a refinement of the original MS-TCN. MS-TCN++ uses a simple two-stage training and a slightly modified convolutional configuration with dilations. We avoid deeper or more nuanced architectures (such as those incorporating attention, or longer context windows) to ensure that performance is due to the quality of the features and not the complexity of the classifier itself.
@@ 4 Experiments
Implementation: We implement multiple Bridge-Prompt variants as well as two baseline image encoders (3DResNet [4] and I3D [2]) in PyTorch. Each Bridge-Prompt implementation is composed of a backbone image encoder (either the default ViT-B/16 [35] or ResNet-50 [36]) and the analogous GPT-2 text encoder [34] which is discarded after training. These backbones are initialized using standard pre-trained weight-sets [^3] Both Bridge-Prompt and 3DResNet methods prescribe the use of weights from another task and training phase before the prompt-tuning phase.. The Bridge-Prompt encoders are then prompt-tuned using the Adam optimizer minimizing the sum of the losses in Eq. 3 in the prompt-tuning phase.
After prompt-tuning we freeze the weights of each image encoder, then train a standard prediction network for the supervised task (MS-TCN++ [12]). This second directly supervised phase we call the supervised training phase. The supervised training phase has its own hold-out test set and it is on this hold-out that we report performance metrics. For JIGSAWS we use Leave-One-User-Out (LOUO) [11] cross-validation, and for RARP-45 we choose 10 videos (out of 36 total) as the test set.
@@
Training time: All experiments were run on NVIDIA A40 GPUs or NVIDIA A5000 GPUs. The prompt-tuning phase for all Bridge-Prompt variants was conducted on two GPUs, otherwise only a single GPU was used. In JIGSAWS, pre-training for 50 epochs for each tested variant of Bridge-Prompt takes approximately 8 hours using two A40 GPUs. It takes 5 minutes to train the MS-TCN++ during the supervised training phase, using pre-extracted image encodings.
Performance metrics: We assess the outcomes using five standard evaluation metrics, Accuracy (Acc.), Edit Distance, and F1@{10,25,50} [24]. For JIGSAWS we condition this on pre-training task (Knot Tying or Suturing), and for zero-shot cases we also condition on each unseen label.
@@
We first measure the quality of the features learned by the various encoders operating under normal (non-zero-shot) conditions. Results are reported in Table 1 and 2. These results show that Bridge-Prompt improves gesture recognition performance as measured by all but one performance metric. Further, due to the similarity between Bridge-Prompt performance with either ResNet50 or ViT backbones, this performance gain does not appear to be due to the transformer architecture of ViT. However, ResNet50 backbones appear to be slightly less stable in RARP-45 training; across multiple runs we experienced exploding gradients, and the method failed to converge. Tuning the learning-rate might resolve this issue, but we did not have the resources to tune that parameter. We include JIGSAWS results reported in van Amsterdam et al. 2022 for contextualization [^4] Van Amsterdam et al. 2022 [6] reports results for models that include input data from both video and robotic kinematic streams, and thus are not entirely comparable for selecting video encoders. Moreover, the particular architecture, training details, and data split cannot be reproduced without access to their codebase. We exclude their RARP-45 reported values as they evaluate on all 45 videos, but only 36 videos in RARP-45 are publicly available., as we believe this to be a state-of-the-art contemporary system utilizing all available data streams (both visual and kinematic), and (presumably) many model architecture optimizations and procedure refinements.
Table 1: Here we present gesture recognition results on the JIGSAWS datatset for 3DResNet, I3D, and Bridge Prompt (BP) with both ResNet50 and ViT variants, as well as the Bridge Prompt architecture with no JIGSAWS specific prompt-tuning (CLIP). We provide the MA-TCN[C] numbers reported by van Amsterdam et al. [6] for overall context, but MA-TCN[C] uses both the video and kinematic data streams and a more complex prediction model.
@@
Table 2: Here we present gesture recognition results on the RARP-45 datatset for 3DResNet, I3D, and Bridge Prompt (BP) with the ViT variant, as well as a “no RARP-45 specific prompt-tuning” case (CLIP). BP-ResNet50 failed to converge, thus we decide not to post it.
Acc. Edit F1@10 F1@25 F1@50
3DResNet 66.97 76.95 71.76 66.34 52.90
I3D 65.95 74.41 71.52 65.35 51.23
CLIP-ViT 70.00 73.88 74.94 73.30 59.95
CLIP-ResNet50 68.93 71.85 70.83 66.90 55.09
BP-ViT 77.36 72.16 76.29 74.50 65.32
[…]
Zero-shot Capability: We measure the efficacy of the Bridge-Prompt image encoder for describing unseen gestures, i.e., zero-shot generalization capability. This is done by selectively training encoders using only subsets of the gestures: for JIGSAWS, we test subsets with only gestures 1-5, only gestures 1-10, and no gestures at all (the “pure CLIP [29] encoder”), presenting results in Table 3. We disaggregate performance by task, gesture, and model in Table 5. The 1-10 gesture model necessarily contains 6 and 8, and thus performance is not reported. For RARP-45 we test subsets with only gestures 1-3 and gestures 1-6, presenting results in Table 4. Per-gesture performance disaggregation is provided as a table in the Appendix. For JIGSAWS, we also trained video encoders on one task but evaluated them on the other (“cross-task”), reported at the bottom of Table 3. This experiment is not possible in RARP-45, as it only has one task. In general we find that Bridge-Prompt prompt-tuning with only a subset of gestures or even on a different task still provides improvement for most performance measures. This, to us, indicates that Bridge-Prompt has relatively high zero-shot capability.
Table 3: Above the double line, we show zero-shot gesture recognition on JIGSAWS with different sets of pre-trained labels (%), along with an All gestures supervised case for reference, and encoders using indices (e.g., “Gesture 2”) instead of text descriptions during prompt-tuning (“ind”). “No gestures” has no prompt-tuning. Below the double line, we show cross-task zero-shot generalization, using the Knot-Tying prompt-tuned model for Suturing supervised training/evaluation, and vice versa.
@@
Table 4: Here we show zero-shot surgical gesture recognition on RARP-45 with different sets of pre-trained labels (%) on RARP-45 dataset, and the same encoders using indices (e.g., “Gesture 2”) instead of text descriptions during prompt-tuning (“ind”).
Acc. Edit F1@10 F1@25 F1@50
3 gestures 77.78 75.09 80.96 79.28 70.60
3 gestures (ind) 78.93 75.80 80.82 79.38 69.78
6 gestures 79.19 71.85 76.60 73.68 65.17
6 gestures (ind) 79.73 76.51 80.70 79.06 69.01
[…]
Table 5: We show zero shot F1@10 scores for single gestures on JIGSAWS with different sizes of pre-training sets (%) that exclude those gestures, along with an All gestures supervised case for reference, and encoders using indices (e.g., “Gesture 2”) instead of text descriptions during prompt-tuning (“ind”).
@@
Text description ablation: Finally, we show evidence that text descriptions provide negligible benefit for Bridge-Prompt encodings. We measured the value of gesture text descriptions by ablating them to simple “Gesture Index” categorical descriptors. For example, gesture 9 could be described as “using right hand to help tighten suture” or as “Gesture 9”. Results from these experiments are included in Table 3, 4, and 5.